What is a situation where there are 12 possible outcomes?
William has 4 shirts that he can wear with a choice of 3 different pants. He has 12 different possible outcomes.
Why there are 12 possible outcomes: There are twelve possible outcomes because 4 multiplied by 3 is 12. This is called algorthim. Algorithm is when you multiply numbers to get your outcome.
There are plenty ways to get or find how many outcomes there are. These ways are a tree diagram, chart, sample spaces, ect.
Hope i helped! :)
How many of the 5 card poker hand possible how many contain at least one card that is not a heart?
There are 2597673 such hands.
There are 2597673 such hands.
How do you remove rusted bolts?
If they are beyond repair, I cut them off with a thin metal cutting disc in an angle grinder.
What is the difference between differential and derivative in the field of calculus?
The derivative refers to the rate at which a function changes with respect to another measure.
The differential refers to the actual change in a function across a parameter.
The differential of a function is equal to its derivative multiplied by the differential of the independent variable .
The derivative of a function is the the LIMIT of the ratio of the increment of a function to the increment of the independent variable as the latter tends to zero.
Probability of an event not occurring?
If an event has a probability of occurring p, then the probability of it not occurring is 1 - p. This is because the probability of something happening is always 1, i.e. p + (1 - p) = 1.
it's 20/50 times 19/49 times 18/48 tmes 17/47 etc down to 1/31
What is non normal distribution?
A distribution that is NOT normal. Most of the time, it refers to skewed distributions.
How do you find standard deviation?
First, you need to determine the mean.
Then, subtract the mean from every number you have. The SQUARE all your numbers.
Add up all of the resulting squares to get their total sum. Divide by one less then the total numbers you have (if you have 6 numbers you will divde by five)
To get the standard deviation, just take the square root of the resulting number
From 0 to 1 or between an event not happening and an event that will happen
simply 10 x 2 so 20
What is the possibility of rolling two dice and getting 2 on both?
The chance of getting a 2 on the first cube is 1 out of 6.
The chance of getting a 2 on the second cube is 1 out of 6.
The chance of getting a 2 on both cubes is (1/6) x (1/6) = 1/36 = 2.78% (rounded)
What if 2 dice are rolled 100 times what is the probability that it will be a prime number?
There are 36 permutations of two standard 6 sided dice. Of those, 15 are prime, namely 2, 3, 3, 5, 5, 5, 5, 7, 7, 7, 7, 7, 7, 11, and 11. The probability of rolling a prime number, then, is 15 in 36, or 5 in 12, or about 0.4167. It does not matter how many times you roll the dice.
A card is drawn at random from a deck what are the odds of drawing a king or a red card?
The probability of drawing a king or a red card from a standard deck of 52 cards is (4 + 26 - 1) in 52, or 29 in 52, or about 0.5577.
There are four kings and 26 red cards, of which one is a king. Simply count the number of desired outcomes (29) and divide by the number of possible outcomes (52).
What are the best numbers to use to win the lottery?
The ones within the range of those to be chosen.
nobodys gonna answer that weird question
If you have a 6 sided dice what is the fraction of getting a 5?
The fraction of getting a 5 is 1/6 because there are 6 sides
If you draw a card from a standard deck of cards what is the probability of drawing a spade?
To determine probability you do
Number of Times the Desired Event will occur
--------------------------------------------------------
Number of Possible Outcomes
In this case, there are 52 cards in a deck (assuming no jokers).
1/4th of those cards are spades (This gives us 13 spades)
so Number of Times the Desired Event will occur = 13
Number of Possible outcomes = 52
You will draw a spade 13 out of 52 times. (13/52)
Or reduced - You will draw a spade 1 out of 4 times (1/4)
The probability of getting 3 white balls in a draw of 5 balls with replacement from an urn containing white balls and black balls is always greater than the same test without replacement, because the number of white balls decreases when you draw a white ball and do not replace it.
The ratio of white to black with replacement is constant, and is always less than one, assuming there is at least one black ball. The ratio of white to black without replacement decreases each turn, and is still less than one, and is less than the previous ratio, unless the question asked about 2 white balls or less.
When finding standard deviation you are typically looking at what kind of distribution?
Usually a normal distribution.
What is the probability of 3 green marbles?
the probability is you'd get a green marble any other color is impossible. So, the probability is certain
The sum of the probability of success and the probability of failure
had better equal 1.00 (100%) or an error has been committed.
The integral of the density function from the given point upwards.
What is the probability of selecting a number less than 4 between 1 to 25?
It is 3/25 or 12%
It is 3/25 or 12%
It is 3/25 or 12%
It is 3/25 or 12%
The probability that an offspring will have wrinkled seeds is 2 in 4 or 50%